express 0.671 ( with a repeating line over the 71) as a fraction.
Answer by user_dude2008(1862)

let x = 0.671717171


---> 1000x = 671.717171


----> 100000x = 67171.7171


Now subtract and solve:

100000x - 1000x = 67171.717171 - 671.717171


99000x = 66500


x = 66500/99000


x = 133/198


Answer: 0.671717171717171... = 133/198